# Linear Complexity Computation of Code Distance and Minimum Size of Trapping Sets for LDPC Codes with Bounded Treewidth

### v1

@@ Abstract
Index Terms:
low-density parity-check codes, tree decomposition, trapping sets, NP-complete.
@@ I Introduction
Instance: Graph $G=(V,E)$, $k,r\in\mathbb{N}$.
@@
Question: Let a $w$-dominating set be a vertex subset $S$ such that for all $v\in V$, either $v\in S$ or it has at least $w$ neighbors in $S$. Decide whether $G$ has a $w$-dominating set of cardinality at most $r$.
@@ II-A Trapping Set
Let $G=(\mathcal{L}\cup\mathcal{R},E)$ be a bipartite graph, or Tanner graph, corresponding to the LDPC code, where $\mathcal{L}$ is the set of variable nodes, $\mathcal{R}$ is the set of check nodes, and $E$ represents the edges between $\mathcal{L}$ and $\mathcal{R}$. For a node $v\in\mathcal{L}\cup\mathcal{R}$, its degree is the number of neighbors of $v$. For subsets $L\subset\mathcal{L}$ and $R\subset\mathcal{R}$, $G[L\cup R]$ denotes the induced subgraph of $L\cup R$ in the Tanner graph. This induced subgraph has $L\cup R$ as the set of vertices, and the set of edges is given by $\{(v,c)|v\in\mathcal{L},c\in\mathcal{R},(v,c)\in E\}$. Let $\mathcal{N}_{G}(L)$ denote the set of neighbors of $L$ in $G$, we can use $G[L]$ as a shorthand for $G[L\cup\mathcal{N}_{G}(L)]$. Next, we provide the definition of a trapping set.
@@
For a Tanner graph $G=(\mathcal{L}\cup\mathcal{R},E)$ and a non-empty subset $S\subset\mathcal{L}$, let $\Gamma_{o}(G[S])$ denote the set of check nodes with odd degrees in the induced subgraph $G[S]$. We define $S$ as an $(a,b)$-trapping set in $G$ if $|S|=a$ and $|\Gamma_{o}(G[S])|=b$. The integer $a$ is referred to as the size of the trapping set $S$.
@@ II-B Treewidth
Tree decomposition and treewidth are important concepts in graph theory. We first provide their definitions for general graphs and then apply them to Tanner graphs.
Let $G=(V,E)$ be a graph, where $V$ is the set of vertices and $E$ is the set of edges. Let $T$ be a tree. The family $(B_{t})_{t\in T}$ is a collection of vertex subsets $B_{t}\subset V$ indexed by the nodes of $T$. Then, tree decomposition can be defined as follows:
@@
The tree decomposition of $G=(V,E)$ is a pair $(T,(B_{t})_{t\in T})$ such that:
@@
    for every edge $(u,v)\in E$, there exists a $t\in T$ such that $u\in B_{t}$ and $v\in B_{t}$;
@@
    for every vertex $v$, the set $B^{-1}(v)=\{t\in T|v\in B_{t}\}$ is non-empty and connected in $T$.
@@
The treewidth of a graph $G$, denoted as $tw(G)$, is the minimum width among all tree decompositions of $G$. The path-width of a graph $G$, denoted as $pw(G)$, is the minimum width among all tree decompositions of $G$ in which $T$ is a path.
As defined in Definition 2, for each graph $G=(V,E)$, there exists a tree decomposition $(T,(B_{t})_{t\in T})$ where $T$ contains only a single node $t$ such that $B_{t}=V$. This implies that the treewidth of $G$ is at most |V|1|V|-1. A rooted tree decomposition $(T,(B_{t})_{t\in T})$ is a tree decomposition where $T$ has a single root node. For example, Fig.1 illustrates two rooted tree decompositions on a bipartite graph.
@@
    for the root node or leaf node $t\in T$, $B_{t}=\emptyset$;
@@
    if a node $t\in T$ has two child nodes $t_{1}$ and $t_{2}$, then $B_{t}=B_{t_{1}}=B_{t_{2}}$. node $t$ is called a join node;
  4. 4.
    if a node $t\in T$ has a single child node $t_{1}$, then one of the following two conditions must hold:
    • There exists a vertex $v\in V$ such that $B_{t}=B_{t_{1}}\cup\{v\}$. In this case, node $t$ is called an introduced node;
    • There exists a vertex $v\in V$ such that $B_{t}=B_{t_{1}}\setminus\{v\}$. In this case, node $t$ is called a forget node.
@@
The above definition and discussion also apply to the case where $G$ is a bipartite graph. In particular, we have the definition as follows:
Definition 4.
Let $G=(\mathcal{L}\cup\mathcal{R},E)$ be a Tanner graph with a tree decomposition $(T,(B_{t})_{t\in T})$.
  1. 1.
    For each bag $B_{t}$ with $t\in T$, we define $B_{t}^{c}$ to denote the set of check nodes in $B_{t}$, and $B_{t}^{v}$ to denote the set of variable nodes in $B_{t}$.
  2. 2.
    For each node $t\in T$, we define $T_{t}$ as the induced subgraph of $T$ consisting of all descendants of $t$ in $T$, and $G_{t}=G[\cup_{i\in T_{t}}B_{i}]$.
  3. 3.
    For each node $t\in T$, we define $\mathcal{L}(G_{t})$ and $\mathcal{R}(G_{t})$ as the sets of variable nodes and check nodes in $G_{t}$, respectively.
From the above definitions, we have $\mathcal{L}(G_{t})=\cup_{i\in T_{t}}B_{i}^{v}$ and $\mathcal{R}(G_{t})=\cup_{i\in T_{t}}B_{i}^{c}$.
@@ II-C Set Operations
  • Power Set $(2^{A})$: The power set of $A$, denoted as $2^{A}$, is the set of all subsets of $A$, including the empty set and $A$ itself:
    $2^{A}=\{X\mid X\subseteq A\}.$
@@
  • $(A\setminus B)\cap(U\setminus W)=(A\cap U)\setminus(B\cap W)$.
@@ III Computer the Size and Number of the Smallest Trapping Set with $b=0$ in Codes with Bounded Treewidth
Consider a bipartite graph $G=(\mathcal{L}\cup\mathcal{R},E)$ corresponding to the LDPC code $\mathcal{C}$, with treewidth $k$. Next, for a given $b$, we will prove that finding an $(a,b)$-trapping set with the minimum $a$ in $G$ can be solved in linear complexity with respect to the code length. In this section, we focus on the case where $b=0$, while the generalization to $b>0$ will be presented in Section IV. The motivation for first considering the case $b=0$ is twofold. On the one hand, the $(a,0)$-trapping set is the most influential structure affecting the error floor of LDPC codes, and finding the smallest trapping set with $b=0$ in $G$ has received significant attention. On the other hand, the case $b>0$ is more complex. To enhance clarity and facilitate understanding, we start with the simpler case of $b=0$. Furthermore, we also show that the number of smallest trapping sets with $b=0$ contained in $G$ can be computed in linear complexity with respect to the code length.
@@ III-A Several Important Properties of Tree Decomposition
To better utilize tree decomposition in designing algorithms for finding the smallest trapping set, we introduce several important properties to deepen our understanding of tree decomposition.
Consider a rooted nice tree decomposition of $G$, denoted by $(T,(B_{t})_{t\in T})$. Let $t$ be a node in $T$. The concepts of $G_{t}$, $B_{t}^{c}$, $B_{t}^{v}$, and other related concepts are provided in Definition 4.
Property 1: Let $TS,TS_{1},TS_{2}$ be three trapping sets in $G_{t}$, where $TS=TS_{1}\cup TS_{2}$ and $TS_{1}\cap TS_{2}=\emptyset$. Then,
$\Gamma_{o}(G_{t}[TS])=\Gamma_{o}(G_{t}[TS_{1}])\oplus\Gamma_{o}(G_{t}[TS_{2}]),$
$\Gamma_{o}(G_{t}[TS_{1}])=\Gamma_{o}(G_{t}[TS])\oplus\Gamma_{o}(G_{t}[TS_{2}]),$
$\Gamma_{o}(G_{t}[TS_{2}])=\Gamma_{o}(G_{t}[TS])\oplus\Gamma_{o}(G_{t}[TS_{1}]).$
Property 2: Let $S$ be a trapping set in $G_{t}$, and let $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are the subset of $\mathcal{R}(G_{t})$ such that $\mathcal{R}_{1}\cup\mathcal{R}_{2}=\mathcal{R}(G_{t})$, and $\mathcal{R}_{1}\cap\mathcal{R}_{2}=\emptyset$. Then,
$\begin{split}\Gamma_{o}(G_{t}[S])&=\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{1}])\cup\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{2}])\\ &=\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{1}])\oplus\Gamma_{o}(G_{t}[S\cup\mathcal{R}_{2}]).\end{split}$
Property 3: Let $L\subset\mathcal{L}(G_{t})$, $R\subset\mathcal{R}(G_{t})$ be two subsets. Then $G_{t}[L\cup R]=G[L\cup R].$
Property 4: Let $t\in T$ be a join node with two child nodes $t^{\prime}$, where $\cup_{i\in T_{t}}B_{i}^{c}=\cup_{i\in T_{t^{\prime}}}B_{i}^{c}$. $S$ is a trapping set in both $G_{t}$ and $G_{t^{\prime}}$. Then,
$\Gamma_{o}(G_{t}[S])=\Gamma_{o}(G_{t^{\prime}}[S]).$
Let $t\in T$ be a join node with two child nodes $t_{1}$ and $t_{2}$, where $B_{t}=B_{t_{1}}=B_{t_{2}}$. Then,
Property 5: $\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\cap\left(\cup_{i\in T_{t_{2}}}B_{i}\right)=B_{t}.$
Property 6: for any nodes $u\in\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\setminus B_{t}$ and $v\in\left(\cup_{i\in T_{t_{2}}}B_{i}\right)\setminus B_{t}$, there cannot exist an edge between $u$ and $v$ in $G$.
Property 7: $\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\setminus B_{t}$, $B_{t}$, and $\left(\cup_{i\in T_{t_{2}}}B_{i}\right)\setminus B_{t}$ are three mutually disjoint sets whose union equals $\cup_{i\in T_{t}}B_{i}$.
The proofs of the first two properties mainly rely on the definition of trapping sets and set operations, while the proof of Property 3 primarily depends on the definition of $G_{t}$. The last three properties can be directly derived from the definition of tree decomposition, and thus, this proof is omitted here. Property 4 can be derived from Property 3 as follows:
$\begin{split}&\Gamma_{o}(G_{t^{\prime}}[S])=\Gamma_{o}(G_{t^{\prime}}[S\cup(\cup_{i\in T_{t^{\prime}}}B_{i}^{c})])\\ &\overset{Prop.3}{=}\Gamma_{o}(G_{t}[S\cup(\cup_{i\in T_{t}}B_{i}^{c})])=\Gamma_{o}(G_{t}[S]).\end{split}$
@@ III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
Definition 5.
Let $(T,(B_{t})_{t\in T})$ be a rooted nice tree decomposition of $G$. Consider a node $t\in T$, node sets $R\subset B_{t}^{c}$, $Q\subset B_{t}^{v}$, and a trapping set $S$ in $G$. We define the parameters of $S$ in $G_{t}$ as $(R,Q)$ in this Section if $S$ satisfies the following properties:
  1. 1.
    $S\subset\mathcal{L}(G_{t})$;
  2. 2.
    $S\cap B_{t}^{v}=Q$;
  3. 3.
    the set of check nodes with an odd degree in $G_{t}[S]$ is exactly $R$.
Lemma 1.
Let $(T,(B_{t})_{t\in T})$ be a rooted nice tree decomposition of $G$, and $t\in T$ be a join node with two child nodes $t_{1}$ and $t_{2}$, where $B_{t}=B_{t_{1}}=B_{t_{2}}$. Suppose $S_{t_{1}}$ and $S_{t_{2}}$ are trapping sets in $G_{t_{1}}$ and $G_{t_{2}}$, respectively. The parameters of $S_{t_{1}}$ in $G_{t_{1}}$ are $(R_{t_{1}},Q)$, while the parameters of $S_{t_{2}}$ in $G_{t_{2}}$ are $(R_{t_{2}},Q)$. If the trapping set $S$ is defined as $S=S_{t_{1}}\cup S_{t_{2}}$, then its corresponding parameters in $G_{t}$ are $(R,Q)$, where
$R=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}]).$
Proof.
First, we compute the second parameter associated with $S$ in $G_{t}$. According to Definition 5, we have $Q=S_{t_{1}}\cap B_{t_{1}}^{v}$, and $Q=S_{t_{2}}\cap B_{t_{2}}^{v}$, thus,
$\begin{split}S\cap B_{t}^{v}&=(S_{t_{1}}\cup S_{t_{2}})\cap B_{t}^{v}\\ &=(S_{t_{1}}\cap B_{t}^{v})\cup(S_{t_{2}}\cap B_{t}^{v})\\ &=(S_{t_{1}}\cap B_{t_{1}}^{v})\cup(S_{t_{2}}\cap B_{t_{2}}^{v})\\ &=Q.\end{split}$
Next, we compute the first parameter associated with $S$ in $G_{t}$. The trapping set $S$ can be regarded as the union of three disjoint sets of variable nodes:
(1) $Q$,
(2) $S_{t_{1}}\setminus Q$, denoted as $S_{1}$,
(3) $S_{t_{2}}\setminus Q$, denoted as $S_{2}$.
Sets $Q$, $S_{1}$, or $S_{2}$ may be empty. Let $\Gamma_{o}(G_{t}[\emptyset])=\emptyset$, we can still proceed with the analysis as follows.
For part (1): The set of check nodes with odd degrees in $G_{t}[Q]$ is given by $\Gamma_{o}(G_{t}[Q])$.
For part (2): $S_{1}$ is a set of variable nodes entirely contained in $\left(\cup_{i\in T_{t_{1}}}B_{i}\right)\setminus B_{t}$. According to Property 6, we can conclude that $S_{1}$ and $\left(\cup_{i\in T_{t_{2}}}B_{i}\right)\setminus B_{t}=\mathcal{R}(G_{t})\setminus\mathcal{R}(G_{t_{1}})$ do not have any edges connecting them. According to Property 2, this implies that $\Gamma_{o}(G_{t}[S_{1}])=\Gamma_{o}(G_{t}[S_{1}\cup\mathcal{R}(G_{t_{1}})])=\Gamma_{o}(G_{t_{1}}[S_{1}])$. Furthermore, since $S_{t_{1}}=S_{1}\cup Q$, then,
$\begin{split}\Gamma_{o}(G_{t}[S_{1}])&=\Gamma_{o}(G_{t_{1}}[S_{1}])\\ &\overset{Prop.1}{=}\Gamma_{o}(G_{t_{1}}[S_{t_{1}}])\oplus\Gamma_{o}(G_{t_{1}}[Q])\\ &=R_{t_{1}}\oplus\Gamma_{o}(G_{t_{1}}[Q]).\end{split}$
For part (3): Similar to part (2), we obtain
$\Gamma_{o}(G_{t}[S_{2}])=R_{t_{2}}\oplus\Gamma_{o}(G_{t_{2}}[Q]).$
Then, according to Property 1, for the union of these three parts, $S$, we have
$\begin{split}\Gamma_{o}(G_{t}[S])&=R_{t_{1}}\oplus R_{t_{2}}\oplus\Gamma_{o}(G_{t_{1}}[Q])\\ &\oplus\Gamma_{o}(G_{t_{2}}[Q])\oplus\Gamma_{o}(G_{t}[Q]),\end{split}$ (1)
and
$\begin{split}&\Gamma_{o}(G_{t_{1}}[Q])\oplus\Gamma_{o}(G_{t_{2}}[Q])\\ &\overset{Prop.2}{=}\Gamma_{o}(G_{t_{1}}[Q\cup\mathcal{R}_{1}])\oplus\Gamma_{o}(G_{t_{1}}[Q\cup B_{t_{1}}^{c}])\\ &\quad\quad\;\;\oplus\Gamma_{o}(G_{t_{2}}[Q\cup\mathcal{R}_{2}])\oplus\Gamma_{o}(G_{t_{2}}[Q\cup B_{t_{2}}^{c}])\\ &\overset{Prop.3}{=}\Gamma_{o}(G_{t}[Q\cup\mathcal{R}_{1}])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\\ &\quad\quad\;\;\oplus\Gamma_{o}(G_{t}[Q\cup\mathcal{R}_{2}])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}])\\ &\overset{(a)}{=}\Gamma_{o}(G_{t}[Q])\oplus\Gamma_{o}(G_{t}[Q\cup B_{t}^{c}]),\end{split}$ (2)
where $\mathcal{R}_{1}=(\cup_{t\in T_{t_{1}}}B_{i}^{c})\setminus B_{t_{1}}^{c}$ and $\mathcal{R}_{2}=(\cup_{t\in T_{t_{2}}}B_{i}^{c})\setminus B_{t_{2}}^{c}$. Equation (a) holds because, according to Property 7, $\mathcal{R}_{1}$, $\mathcal{R}_{2}$, and $B_{t}^{c}$ are three mutually disjoint sets whose union forms $\cup_{i\in T_{t}}B_{i}^{c}=\mathcal{R}(G_{t})$. This result follows from Property 2.
Combining Equations (1) and (2) completes the proof of the Lemma.
@@
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. Then, the size and the number of the smallest trapping sets with $b=0$ in $G$ can be determined in time $O(n2^{2k})$.
@@
A nice tree decomposition of $G$ is given by $(T,(B_{t})_{t\in T})$, with root $r$ and width $k$. For each bag $B_{t}$, we define two functions,
$f_{t}:(L\in 2^{B_{t}^{c}},Q\in 2^{B_{t}^{v}})\mapsto f_{t}(L,Q)\in\mathbb{Z},$
@@
$g_{t}:(L\in 2^{B_{t}^{c}},Q\in 2^{B_{t}^{v}})\mapsto g_{t}(L,Q)\in\mathbb{Z}.$
where $f_{t}(R,Q)$ denotes the size of the smallest trapping set in $G_{t}$ with parameters $(R,Q)$, while $g_{t}(R,Q)$ represents the number of such trapping sets. Note that $\emptyset$ is not a trapping set. Therefore, even if $\emptyset\cap B_{t}^{v}=\emptyset$ and $\Gamma_{o}(G_{t}[\emptyset])=\emptyset$, we don’t have $f_{t}(\emptyset,\emptyset)=0$ and $f_{t}(\emptyset,\emptyset)=1$.
Since all and only the trapping sets with $b=0$ correspond to the parameters $(\emptyset,\emptyset)$ in $G_{r}$, it follows that once we obtain $f_{r}(\emptyset,\emptyset)$ and $g_{r}(\emptyset,\emptyset)$, we can determine the size and the number of the smallest trapping sets with $b=0$ in the Tanner graph $G$.
Leaf node: Suppose $t$ is a leaf node. Then,
$f_{t}(R,Q)=+\infty,$
@@
$g_{t}(R,Q)=0,$
for all $R\in 2^{B_{t}^{c}}=\emptyset$ and $Q\in 2^{B_{t}^{v}}=\emptyset$.
It takes $O(1)$ time to compute all $f_{t}(R,Q)$ and $g_{t}(R,Q)$ for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$ for each leaf node $t$. After computing $f_{t}$ and $g_{t}$ for the leaf nodes, we then visit the bags of the tree decomposition from the leaves to the root, calculating the corresponding $f$-functions and $g$-functions at each step according to the following rules.
Let $t$ be a non-leaf node and assume we have count $f_{t^{\prime}}(R,Q)$ and $g_{t^{\prime}}(R,Q)$ , for all $R\in 2^{B_{t^{\prime}}^{c}}$ and $Q\in 2^{B_{t^{\prime}}^{v}}$, where $t^{\prime}$ is a child node of $t$. We then design an algorithm for five types of $t$ to compute $f_{t}(R,Q)$ and $g_{t}(R,Q)$, for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$. We only provide a justification for the computation of $f_{t}(R,Q)$, as the validity of $g_{t}(R,Q)$ follows directly from the proof of $f_{t}(R,Q)$.
introduced variable node: Suppose $t$ is an introduced node, and $t^{\prime}$ is its only child node, with $B_{t}=B_{t^{\prime}}\cup\{v\}$, where $v$ is a variable node. This means the functions $f_{t}$ and $g_{t}$ are extensions of the domains of $f_{t^{\prime}}$ and $g_{t^{\prime}}$, respectively, as $(2^{B_{t}^{c}}=2^{B_{t^{\prime}}^{c}},2^{B_{t}^{v}}=2^{B_{t^{\prime}}^{v}\cup\{v\}})$. We can compute $f_{t}$ and $g_{t}$ using the equations (3) and (4) for all $R\in 2^{B_{t}^{c}}$ and $Q\in 2^{B_{t}^{v}}$, where $\Gamma_{o}(G_{t}[v])=\Gamma_{o}(G_{t}[v\cup B_{t}^{c}])$.
@@
$f_{t}(R,Q)=\begin{cases}f_{t^{\prime}}(R,Q),&\text{if }v\notin Q;\\ 1,&\text{if }Q=\{v\}\text{and }R=\Gamma_{o}(G_{t}[v]);\\ f_{t^{\prime}}(R\oplus\Gamma_{o}(G_{t}[v]),Q\backslash\{v\})+1,&\text{otherwise.}\end{cases}$ (3)

### v2

@@ I Introduction
Instance: Graph $G=(\mathcal{V},\mathcal{E})$, $k,r\in\mathbb{N}$.
@@
Question: Let a $w$-dominating set be a vertex subset $S$ such that for all $v\in\mathcal{V}$, either $v\in S$ or it has at least $w$ neighbors in $S$. Decide whether $G$ has a $w$-dominating set of cardinality at most $r$.
@@ II-A Trapping Set
Let $G=(\mathcal{L}\cup\mathcal{R},\mathcal{E})$ be a bipartite graph, or Tanner graph, corresponding to the LDPC code, where $\mathcal{L}$ is the set of variable nodes, $\mathcal{R}$ is the set of check nodes, and $\mathcal{E}$ represents the edges between $\mathcal{L}$ and $\mathcal{R}$. For a node $u\in\mathcal{L}\cup\mathcal{R}$, its degree is the number of neighbors of $u$. For a subset $S\subset\mathcal{L}$, let $\Gamma_{G}(S)$ denote the set of neighbors of $S$ in $G$. The induced subgraph of $S\cup\Gamma_{G}(S)$ in the Tanner graph, denoted $G[S\cup\Gamma_{G}(S)]$, has $S\cup\Gamma_{G}(S)$ as the set of vertices, and the set of edges is given by $\{(v,c)|v\in S,c\in\Gamma_{G}(S),(v,c)\in\mathcal{E}\}$. Let $\Gamma_{G}^{o}(S)$ denote the set of check nodes in $\Gamma_{G}(S)$ with odd degree in $G([S\cup\Gamma_{G}(S)])$. For notational simplicity, throughout this paper we also use $G[S]$ to denote $G[S\cup\Gamma_{G}(S)]$.
@@
For a Tanner graph $G=(\mathcal{L}\cup\mathcal{R},\mathcal{E})$ and a non-empty subset $S\subset\mathcal{L}$, we define $S$ as an $(a,b)$ trapping set in $G$ if $|S|=a$ and $|\Gamma_{G}^{o}(S)|=b$. The integer $a$ is referred to as the size of the trapping set $S$.
@@ II-B Treewidth
Tree decomposition and treewidth are important concepts in graph theory [34]. LDPC codes with bounded treewidth are a class of codes that this paper focuses on.
Consider a bipartite graph $G=(\mathcal{L}\cup\mathcal{R},E)$. Let $T$ be a tree, and $(B_{t})_{t\in T}$ be a collection of node subsets $B_{t}\subset\mathcal{L}\cup\mathcal{R}$ indexed by the nodes of $T$. Then, the tree decomposition can be defined as follows:
@@
The tree decomposition of $G=(\mathcal{L}\cup\mathcal{R},\mathcal{E})$ is a pair $(T,(B_{t})_{t\in T})$ such that:
@@
    for every edge $(v,c)\in\mathcal{E}$, there exists a $t\in T$ such that $v,c\in B_{t}$;
@@
    for every node $u\in\mathcal{L}\cup\mathcal{R}$, the set $B^{-1}(u)=\{t\in T|u\in B_{t}\}$ is non-empty and connected in $T$.
@@
The treewidth of a graph $G$, denoted as $tw(G)$, is the minimum width among all tree decompositions of $G$. The pathwidth of a graph $G$, denoted as $pw(G)$, is the minimum width among all tree decompositions of $G$ in which $T$ is a path.
As defined in Definition 2, for each graph $G=(\mathcal{L}\cup\mathcal{R},\mathcal{E})$, there exists a tree decomposition $(T,(B_{t})_{t\in T})$ where $T$ contains only a single node $t$ such that $B_{t}=\mathcal{L}\cup\mathcal{R}$. This implies that the treewidth of $G$ is at most $|\mathcal{L}\cup\mathcal{R}|-1$. A rooted tree decomposition is a tree decomposition with a distinguished root node. For example, Fig.1 illustrates two rooted tree decompositions on a bipartite graph.
@@
    if a node $t\in T$ has two child nodes $t^{\prime}$ and $t^{\prime\prime}$, then $B_{t}=B_{t^{\prime}}=B_{t^{\prime\prime}}$. node $t$ is called a join node;
@@
    if a node $t\in T$ has a single child node $t^{\prime}$, then one of the following two conditions must hold:
    • There exists a vertex $u\in\mathcal{L}\cup\mathcal{R}$ such that $B_{t}=B_{t^{\prime}}\cup\{u\}$. In this case, node $t$ is called an introduced node;
    • There exists a vertex $u\in\mathcal{L}\cup\mathcal{R}$ such that $B_{t}=B_{t^{\prime}}\setminus\{u\}$. In this case, node $t$ is called a forget node.
@@ II-C Set Operations
  • $(A\setminus B)\cap(U\setminus W)=(A\cap U)\setminus(B\cap W)$,
  • $A\setminus(B\setminus C)=(A\setminus B)\cup(A\cap C)$
  • $(A\oplus B)\setminus C=(A\setminus(B\cup C))\cup(B\setminus(A\cup C))$
  • $A\setminus(B\cup C)=(A\setminus B)\setminus C$
@@ III Computer the Size and Number of the Smallest Trapping Set with $b=0$ in Codes with Bounded Treewidth
Consider a bipartite graph $G=(\mathcal{L}\cup\mathcal{R},\mathcal{E})$ corresponding to the LDPC code $\mathcal{C}$, with treewidth $k$. Next, for a given $b$, we will prove that finding an $(a,b)$-trapping set with the minimum $a$ in $G$ can be solved in linear complexity with respect to the code length. In this section, we focus on the case where $b=0$, while the generalization to $b>0$ will be presented in Section IV. The motivation for first considering the case $b=0$ is twofold. On the one hand, the $(a,0)$-trapping set is the influential structure affecting the error floor of LDPC codes, and finding the smallest trapping set with $b=0$ in $G$ has received significant attention. On the other hand, the case $b>0$ is more complex. To enhance clarity and facilitate understanding, we start with the simpler case of $b=0$. Furthermore, we also show that the number of smallest trapping sets with $b=0$ contained in $G$ can be computed in linear complexity with respect to the code length.
@@ III-A Several Important Properties of Tree Decomposition
To better design algorithms for finding the smallest trapping set, we introduce several important properties to deepen our understanding of trapping set.
Property 1: Let $S,S_{1},S_{2}$ be three trapping sets in $G$, where $S=S_{1}\cup S_{2}$ and $S_{1}\cap S_{2}=\emptyset$. Then,
$\Gamma^{o}_{G}(S)=\Gamma^{o}_{G}(S_{1})\oplus\Gamma^{o}_{G}(S_{2}).$
and
$\Gamma^{o}_{G}(S_{1})=\Gamma^{o}_{G}(S)\oplus\Gamma^{o}_{G}(S_{2}).$
Property 2: Let $S$ be a trapping set in $G$, and let $\mathcal{L}^{\prime}$ be a check node sets in $G$.
$\Gamma^{o}_{G}(S)=\Gamma^{o}_{G\setminus\mathcal{L}^{\prime}}(S)\oplus\Gamma^{o}_{G[\mathcal{L}^{\prime}\cup S]}(S).$
@@ III-B Compute The Size and Number of The Smallest Trapping Set with $b=0$
Let $G$ be a Tanner graph with treewidth $tw(G)=k$, corresponding to an LDPC code of length $n$. Then, for given $b=0$, the size and the number of the smallest $(a,0)$ trapping sets in $G$ can be determined in time $O(n4^{k})$.
@@
A nice tree decomposition of $G$ is given by $(T,(B_{t})_{t\in T})$ with width $k$. Without loss of generality, we may assume that each bag $B_{t}$ corresponding to a leaf node or the root node in $T$ is an empty set. Based on this decomposition, we develop a dynamic programming algorithm that solves the MINIMUM TRAPPING SET problem.
The algorithm processes the nodes of $T$ in a bottom-up order $t_{1},t_{2},\ldots,t_{r}$, such that every child of a node $t_{i}$ appears earlier in the order. At each node $t_{i}$, the algorithm maintains information about trapping sets contained in the subgraph induced by the union of the bags processed so far, namely $G_{i}:=G[\cup_{l=1:i}B_{t_{l}}]$. Let $\mathcal{L}_{i}$ denote the set of variable nodes in $G_{i}$, and let $\mathcal{L}^{B}_{i}$ and $\mathcal{R}^{B}_{i}$ denote the the sets of variable nodes and check nodes in $B_{t_{i}}$, respectively.
Definition 4.
Let $S$ be a trapping set in $G$, $L$ a set of variable nodes, $I$ a set of check nodes. We say that $S$ is an $(L,I,t_{i})$-type trapping set if the following conditions hold:
  1. 1.
    $S\subset\mathcal{L}_{i}$,
  2. 2.
    $S\cap B_{t_{i}}=L$,
  3. 3.
    $\Gamma^{o}_{G_{i}}(S)=I$,
Formally, for each node $t_{i}\in T$, we define two functions as follows,
$f_{t_{i}}:(L,I)\mapsto f_{t_{i}}(L,I)\in\mathbb{Z},$
@@
$g_{t_{i}}:(L,I)\mapsto g_{t_{i}}(L,I)\in\mathbb{Z}.$
where $L\subset\mathcal{L}^{B}_{i}$ is a set of variable nodes in $B_{t_{i}}$, $I\subset\mathcal{R}^{B}_{i}$ is a set of check nodes in $B_{t_{i}}$. The value $f_{t_{i}}(L,I)$ denotes the minimum size of a $(L,I,t_{i})$-type trapping set, while $g_{t_{i}}(L,I)$ counts the number of such minimum-size trapping sets.
The value of $f_{t_{i}}$ and $g_{t_{i}}$ are updated according to the type of node $t_{i}$ in the nice tree decomposition, using the values computed at its children. After processing root node $t_{r}$, $f_{t_{r}}(\emptyset,\emptyset)$ gives the minimum size of an $(a,b)$ trapping set contained in $G$, while $g_{t_{r}}(\emptyset,\emptyset)$ gives the number of minimum $(a,b)$ trapping set contained in $G$. The detailed update rules and complexity analysis are provided in the sequel.
Leaf node: Suppose $t_{i}$ is a leaf node in $T$. Then,
$f_{t_{i}}(L,I)=+\infty,$
@@
$g_{t_{i}}(L,I)=0,$
for $L=\emptyset$, $I=\emptyset$. It takes $O(1)$ time to compute all $f_{t_{i}}(L,I)$ and $g_{t_{i}}(L,I)$ for each leaf node $t_{i}$.
introduced variable node: Suppose $t_{i}$ is an introduced node with a unique child $t_{j}$, and that the corresponding bags satisfy $B_{t_{i}}=B_{t_{j}}\cup\{v\}$, where $v$ is a variable node. We can compute $f_{t_{i}}$ and $g_{t_{i}}$ using the equations (1) and (2) for all variable node set $L\subset{\mathcal{L}^{B}_{i}}$, and check nodes set $I\subset{\mathcal{R}^{B}_{i}}$.
@@
$f_{t_{i}}(L,I)=\begin{cases}f_{t_{j}}(L,I),&\text{if }v\notin L;\\ 1,&\text{if }L=\{v\},\text{ and }I=\Gamma_{G_{i}}^{o}(v);\\ f_{t_{j}}(L\backslash\{v\},I\oplus\Gamma_{G_{i}}^{o}(v))+1,&\text{otherwise.}\end{cases}$ (1)